OpenStudy (anonymous):

\[Evaluate \int\limits [(2x+1)(x ^{2}+x+9)]\]

OpenStudy (anonymous):

well just expand it..and integrate it one by one :)

OpenStudy (angela210793):

i know how to solve easy integrals :P

OpenStudy (anonymous):

Lol, here's a nice one to think about @angela210793 \[ \int \frac{1}{ln(x-1)}\]

OpenStudy (lgbasallote):

@Omniscience there's an easier way. let u = x^2 + x + 9 du = 2x + 1 therefore the integral becomes \[\int udu\]

OpenStudy (anonymous):

keep in mind if the asker knows u-sub first :)

OpenStudy (angela210793):

i said easyyyy :P :(

OpenStudy (lgbasallote):

who doesnt heh

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

well if he/she does he wont be asking these type of questions :)

OpenStudy (lgbasallote):

and im pretty sure @Beatles knows u-sub

OpenStudy (lgbasallote):

he even knows integration by parts, trig sub and partial fractions

OpenStudy (lgbasallote):

that's why i proposed u-sub ;)

OpenStudy (anonymous):

well i didnt know that. if he knows that much why would he ask this question? :)

OpenStudy (anonymous):

WATEVER.

OpenStudy (lgbasallote):

haha